OpenStudy (anonymous):

Choose the equation of the line passing through the point (1, -5) and perpendicular to y = onefifth x - 10 a. y = 5x + 10 b. y = 5x c. y = -5x - 10 d. y = -5x

OpenStudy (anonymous):

i need help on how to figure this out on my own. i have a dozen other ones like this

OpenStudy (anonymous):

the product of slopes of perpendicular lines =-1 y=(1/5)x-10 so the slope of perpendicular line : -5 y-y1=m(x-x1) y+5=-5(x-1) y+5=-5x+5 y=-5x+5-5 y=-5x

OpenStudy (anonymous):

the answer is D.

OpenStudy (anonymous):

And if its a parallel line?

OpenStudy (anonymous):

has same slope.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

welcome ^__^.